OpenStudy (anonymous):

A roasted turkey is taken from an oven when its temperature has reached 185°F and is placed on a table in a room where the temperature is 64°F. (A)If the temperature of the turkey is 150°F after half an hour, what is its temperature after 45 min?

OpenStudy (anonymous):

(b) When will the turkey cool to 100°F?

OpenStudy (anonymous):

we can do this, but it is going to take a minute

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

first of all you need to work not with the temperatures, but with the difference of the temperatures the initial difference is \(185-64=121\) and the difference after 30 minutes is \(150-64=86\)

OpenStudy (anonymous):

are you supposed to model this as \(A_0e^{kt}\) or can we do it an easier way?

OpenStudy (anonymous):

snap way is to say the temperature difference is \[121\times \left(\frac{86}{121}\right)^{\frac{t}{30}}\] and then compute \[121\times \left(\frac{86}{121}\right)^{\frac{45}{30}}\] \

OpenStudy (anonymous):

or \[121\times \left(\frac{86}{121}\right)^{\frac{2}{3}}\] and then add \(64\)

OpenStudy (anonymous):

im just lost with all of it cuz i keep getting 65 and 247. i know thats not the answer.

OpenStudy (anonymous):

i get 85.72 http://www.wolframalpha.com/input/?i=121*%2886%2F121%29^3%2F2+%2B64

OpenStudy (anonymous):

are you supposed to model this as \[A_0e^{kt}\] we can do it that way too

OpenStudy (anonymous):

thats not it

OpenStudy (anonymous):

Is it 25, if my answer is correct, I will try to help you solve it

OpenStudy (anonymous):

is that fr part b?

OpenStudy (anonymous):

No, a

OpenStudy (anonymous):

I can't see the link, it says I have to log in

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

here its in the attachments.

OpenStudy (anonymous):

I can't help :( I didn't think it would be so complicated,@Directrix , @electrokid , could you guys please help @mel11 ? Thanks :)

OpenStudy (anonymous):

ok well thank you for everything

OpenStudy (anonymous):

Your welcome, hopefully, I can find someone to help you

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

well i just got A i had to plug it into a different math website. thank you so much though. i have been trying to work this out for over an hour.

OpenStudy (anonymous):

thank you @satellite73 and @hedyeh99

OpenStudy (anonymous):

so, did you get the answer?

OpenStudy (anonymous):

yes i did. It took me an hour to figure it out.